A conductivity cell dipped in 0.0005M Nacl gives at 250c a resistance of 13710Ω If the electrodes of the conductivity cell are 0.7cm apart and the area of cross section of the electrode is 0.82cm2,What is the molar conductivity of the solution at 250c?
1
Expert's answer
2020-04-27T01:51:53-0400

k = 1/R x l/a = 1/13710 Ohm x 0.7 cm/0.82 cm2 = 6.23*10-5 Ohm-1cm-1

Molar conductivity = k x 1000 / M = 6.23*10-5 Ohm-1cm-1 x 1000 / 0.0005 mol/dm3 = 124.6 Ohm-1cm2mol-1